# Scaling White-Box Transformers for Vision

### v2

@@ 1 Introduction
Figure 1: (Left) We demonstrate how modifications to the components enhance the performance of the crate model on ImageNet-1K. Details are provided in Section 3. (Right). We compare the FLOPs and accuracy on ImageNet-1K of our methods with ViT [8] and CRATE [41]. CRATE is trained only on ImageNet-1K, while crate-$\alpha$ and ViT are pre-trained on ImageNet-21K. The values of crate-$\alpha$ model correspond to those presented in Table 1.
To this end, this paper provides the first exploration of training crate at different scales for vision, i.e., Tiny, Small, Base, Large, Huge. Detailed model specifications are given in Table 4 of Appendix A. To achieve effective scaling, we make two key changes. First, we identify the vanilla ISTA block within crate as a limiting factor that hinders further scaling. To overcome this, we significantly expand the channels, decouple the association matrix, and add a residual connection, resulting in a new model variant — crate-$\alpha$. It is worth noting that this architecture change still preserves the mathematical interpretability of the model. Second, we propose an improved training recipe, inspired by previous work [33, 41, 34], for better coping the training with our new crate-$\alpha$ architecture.
We provide extensive experiments supporting the effective scaling of our crate-$\alpha$ models. For example, we scale the crate-$\alpha$ model from Base to Large size for supervised image classification on ImageNet-21K [5], achieving 85.1% top-1 accuracy on ImageNet-1K at the Large model size. We further scale the model size from Large to Huge, utilizing vision-language pre-training with contrastive learning on DataComp1B [9], and achieve a zero-shot top-1 accuracy of 72.3% on ImageNet-1K at the Huge model size. [^1] Model configurations are detailed in Table 4 (in Appendix A). These results demonstrate the strong scalability of the crate-$\alpha$ model, shedding light on scaling up mathematically interpretable models for future work.
@@ 3 CRATE-$\alpha$ Model
$\texttt{Prox}_{\eta,\lambda}[\bm{A};\bm{D},\bm{Z}]=\operatorname{ReLU}(\bm{A}-\eta\bm{D}^{\top}(\bm{D}\bm{Z}-\bm{A})-\eta\lambda\bm{1}).$ (9)
@@ 4 Experiments
To thoroughly explore the scalability of crate-$\alpha$ model, we investigate its scaling behaviors from Base to Large size, and ultimately to Huge size. In the previous work [41], the largest model provided is Large size. Thus, for scaling to the Huge model, specifically our crate-$\alpha$-H configuration, we mainly follow the settings of the ViT [8]. For the transition from Base to Large size, we pre-train our model on ImageNet-21K and fine-tune it on ImageNet-1K via supervised learning. When scaling from Large to Huge, we utilize the DataComp1B [9] dataset within a vision-language pre-training paradigm. This allows us to study the effects of scaling the model to a massive size.
4.1 Scaling the crate-$\alpha$ Model from Base to Large
Dataset and evaluation. The standard ImageNet-21K dataset encompasses 21k classes and 14 million images. However, due to data loss, our version of the ImageNet-21K dataset is comprised of 19k classes and 13 million images. We believe that this minor discrepancy does not significantly affect the experimental results. Additionally, during pre-training on ImageNet-21K, we randomly allocated 1% of the data as a validation set to timely monitor the pre-training process. After pre-trained on ImageNet-21K, we further fine-tune the models on ImageNet-1K. We then evaluate the performance of these models on the ImageNet-1K validation set.
@@ 4.1 Scaling the crate-$\alpha$ Model from Base to Large
Training & fine-tuning. We initially pre-train the crate-$\alpha$ model on ImageNet-21K and subsequently fine-tune it on ImageNet-1K. During the pre-training phase, we set the learning rate to $8\times 10^{-4}$, weight decay to 0.1, and batch size to 4096. We apply data augmentation techniques such as Inception crop [30] resized to 224 and random horizontal flipping. In the fine-tuning phase, we adjust the base learning rate to $1.6\times 10^{-4}$, maintain weight decay at 0.1, and batch size at 4096. We apply label smoothing with a smoothing parameter of 0.1 and apply data augmentation methods including Inception crop, random horizontal flipping, and random augmentation with two transformations (magnitude of 9). For evaluation, we resize the smaller side of an image to 256 while maintaining the original aspect ratio and then crop the central portion to 224$\times$224. In both the pre-training and fine-tuning phases, we use the AdamW optimizer [21] and incorporate a warm-up strategy, characterized by a linear increase over 10 epochs. Both the pre-training and fine-tuning are conducted for a total of 91 epochs, utilizing a cosine decay schedule.
Results and analysis. As shown in Table  1, we compare crate-$\alpha$-B and crate-$\alpha$-L at patch sizes 32, 16, and 8. Firstly, we find our proposed crate-$\alpha$-L consistently achieves significant improvements across all patch sizes. Secondly, compared with the results of the vanilla crate (the first row of Table  1), increasing from crate-B to crate-L results in only a 0.5% improvement on ImageNet-1K. This indicates a case of diminishing returns. These findings compellingly highlight that the scalability of crate-$\alpha$ models significantly outperforms that of the vanilla crate. Meanwhile, the training loss in the pre-training stage is presented in Fig. 3; as the model capacity increases, the trend of the training loss improves predictably. This phenomenon is also described in [8].
@@
Models (Base) ImageNet-1K(%) Models (Large) ImageNet-1K(%)
@@
4.2 Scaling the crate-$\alpha$ Model from Large to Huge
Dataset and evaluation. In this part, we further increase the scale of the crate-$\alpha$ model to a Huge size. Building upon the insights from [8], the data-intensive nature of transformers poses significant challenges when attempting to scale vision transformers to a massive model size on ImageNet-21K. Thus, [8] utilizes the JFT-300M dataset to further scale the size of the training data. Since JFT-300M is not publicly accessible, we utilize a larger open dataset, DataComp1B [9]. This new state-of-the-art multimodal dataset, containing 1.4 billion image-text pairs, allows us to further scale our model to a massive size. Specifically, we train crate-$\alpha$ models using a contrastive learning-based approach, following the CLIP [26] paradigm. This enables us to observe scaling benefits when increasing the model size from Large to Huge. Considering the substantial computational requirements of training vanilla CLIP [26], we train the CLIP-like model following the CLIPA [15] protocol, which significantly reduces computational resources while maintaining competitive performance relative to CLIP. For evaluation, we evaluate the zero-shot accuracy of these models on ImageNet-1K.
Training & fine-tuning. In the pre-training stage, we utilize an image size of 84$\times$84, and the maximum token length is 32, with a total of 2.56 billion training samples. During the fine-tuning stage, we increase the image size to 224$\times$224 while maintaining the maximum token length at 32, with a 512 million training samples. Here, the key distinction between the pre-training stage and the fine-tuning stage is the image size. A smaller image size results in a faster training speed. In the configurations of crate-$\alpha$-CLIPA-B, crate-$\alpha$-CLIPA-L, and crate-$\alpha$-CLIPA-H, we use the crate-$\alpha$ model as the vision encoder, and utilize the same pre-trained huge transformer model from CLIPA [15] as the text encoder. For both the pre-training and fine-tuning stages, we freeze the text encoder and only train the vision encoder, i.e., the crate-$\alpha$ model. As we will show in the later results, this setup effectively demonstrates the scaling behaviors of crate-$\alpha$ models in the image domain. Detailed hyperparameter settings can be found in Appendix A.
Results and analysis. From the results shown in Fig. 4, we find that: (1) crate-$\alpha$-CLIPA-L/14 significantly outperforms crate-$\alpha$-CLIPA-B/16 by 11.3% and 9.0% in terms of ImageNet-1K zero-shot accuracy during the pre-training and fine-tuning stages, respectively. The substantial benefit suggests that the quality of learned representation may be constrained by the model size. Therefore, increasing the model size effectively leverages larger amounts of data. (2) When continuing to scale up model size, we also observe that crate-$\alpha$-CLIP-H/14 continues to benefit from larger training datasets, outperforming crate-$\alpha$-CLIP-L/14 by 3.1% and 2.5% in terms of ImageNet-1K zero-shot accuracy during the pre-training and fine-tuning stages, respectively. This demonstrates the strong scalability of the crate-$\alpha$ model. To explore the performance ceiling, we train a standard ViT-CLIPA-H/14 from scratch and observe improved performance.
@@ 4.3 Compute-efficient Scaling Strategy
We further explore methods to scale models efficiently in terms of computation. Table 1 demonstrates that the crate-$\alpha$ model scales effectively from the Base model to its larger variants. However, the pre-training computation for the top-performing model, crate-$\alpha$-L/8, is resource-intensive on ImageNet-21K. Inspired by CLIPA [15], we aim to reduce computational demands by using reduced image token sequence lengths during the pre-training stage, while maintaining the same training setup during the fine-tuning stage. The results are summarized in Table 2.
Results and analysis. (1) When fine-tuning with crate-$\alpha$-L/14 and using crate-$\alpha$-L/32 for pre-training on ImageNet-21K, this approach consumes less than 30% of the TPU v3 core-hours required by crate-$\alpha$-L/14, yet achieves a promising 83.7% top-1 accuracy on ImageNet-1K, comparable to the 83.9% achieved by crate-$\alpha$-L/14; (2) When fine-tuning with crate-$\alpha$-L/8 and using crate-$\alpha$-L/32 for pre-training, this approach consumes just 10% of the training time required by crate-$\alpha$-L/8, yet it still achieves a promising 84.2% top-1 accuracy on ImageNet-1K, compared to 85.1% when using the crate-$\alpha$-L/8 model in the pre-training stage. In summary, we find that this strategy offers a valuable reference for efficiently scaling crate-$\alpha$ models in the future.
Table 2: Compute-efficient scaling strategy. To reduce the compute requirements of the pre-training stage, we use a model with a larger patch size. This results in a shorter token length for the same input size. The second column indicates the compute requirements of the pre-training stage, measured in TPU v3 core-hours. Details are provided in Section 4.3.
Model used in pre-train stage TPU v3 core-hours Fine-tuning model ImageNet-1K(%)
crate-$\alpha$-L/32 2,652 crate-$\alpha$-L/14 83.7
crate-$\alpha$-L/8 84.2
crate-$\alpha$-L/14 8,947 crate-$\alpha$-L/14 83.9
crate-$\alpha$-L/8 35,511 crate-$\alpha$-L/8 85.1
4.4 Improved Semantic Interpretability of crate-$\alpha$
As shown in previous subsections, crate-$\alpha$ can be effectively scaled up from Base to Huge when increasing the data size from millions to billions of images. In addition to scalability, we study the interpretability of our proposed crate-$\alpha$ with different model sizes. we follow the evaluation setup of crate in [41]. Specifically, we apply MaskCut [38] to validate and evaluate the rich semantic information captured by our model, including both qualitative and quantitative results.
As shown in Figure 5, we provide the segmentation visualization on COCO val2017 [17] for crate-$\alpha$, crate, and ViT, respectively. We find that our model preserves and even improves the (semantic) interpretability advantages of crate. Moreover, we summarize quantitative evaluation results on COCO val2017 in Table 3. Interestingly, when scaling up model size for crate-$\alpha$, the Large model improves over the Base model in terms of object detection and segmentation.
@@ 5 Discussion
Societal impact. A possible broader implication of this research is the energy consumption needed to conduct the experiments in our scaling study. But we believe understanding the scaling behaviors of white-box transformers is meaningful for practical purposes due to their mathematical and semantic interpretability. Additionally, our trained models can be used in various downstream tasks in the transfer learning setting, further amortizing the pre-training compute costs.
@@ Appendix A Additional Experiments and Details
Model configuration. We provide details about crate-$\alpha$ model configurations in Table 4.
@@
Training details of crate-$\alpha$-CLIPA models. When employing the crate-$\alpha$ architecture to replace the vision encoder in the CLIPA [15] framework, we essentially follow the original CLIPA training recipe. The setup for the pre-training stage is presented in Table 5. During the fine-tuning stage, we made some modifications: the input image size is set to $224\times 224$, the warmup steps are set to 800, and the base learning rate is set to 4e-7. When calculating the loss, we use the classification token from the vision encoder as the image feature and the last token from the text encoder as the text feature.
@@
Visualization of loss curves. We visualize the training loss curves of the four models, including crate and its three variants, in Fig. 7. We visualize the training loss curves of crate-$\alpha$-Base with different patch sizes in Fig. 8. In Fig. 9, we also visualize the training loss curves of models trained with efficient scaling strategy described in Section 4.3.
Figure 7: Training loss curves of different model architectures (mentioned in Fig. 1) on ImageNet-21K. The patch size is 32 for all four models shown in this figure. (+O: +overcomplete dictionary, +D: +decoupled dictionary, +R: +residual connection.)

### v4

@@ 1 Introduction
Figure 1: (Left) We demonstrate how modifications to the components enhance the performance of the crate model. The four models are trained using the same setup: first pre-trained on ImageNet-21K and then fine-tuned on ImageNet-1K. Details are provided in Section 3. (Right). We compare the FLOPs and accuracy on ImageNet-1K of our methods with ViT [9] and CRATE [46]. The values of crate-$\alpha$ model correspond to those presented in Table 1. A more detailed comparison between crate-$\alpha$ and ViT is included in Appendix A.2.
To this end, this paper provides the first exploration of training crate at different scales for vision, i.e., Tiny, Small, Base, Large, Huge. Detailed model specifications are given in Table 7 of Appendix A.1. To achieve effective scaling, we make two key changes. First, we identify the vanilla ISTA block within crate as a limiting factor that hinders further scaling. To overcome this, we significantly expand the channels, decouple the association matrix, and add a residual connection, resulting in a new model variant — crate-$\alpha$. It is worth noting that this architecture change still preserves the mathematical interpretability of the model. Second, we propose an improved training recipe, inspired by previous work [38, 46, 39], for better coping the training with our new crate-$\alpha$ architecture.
We provide extensive experiments supporting the effective scaling of our crate-$\alpha$ models. For example, we scale the crate-$\alpha$ model from Base to Large size for supervised image classification on ImageNet-21K [6], achieving 85.1% top-1 accuracy on ImageNet-1K at the Large model size. We further scale the model size from Large to Huge, utilizing vision-language pre-training with contrastive learning on DataComp1B [10], and achieve a zero-shot top-1 accuracy of 72.3% on ImageNet-1K at the Huge model size. [^1] Model configurations are detailed in Table 7 (in Section A.1). These results demonstrate the strong scalability of the crate-$\alpha$ model, shedding light on scaling up mathematically interpretable models for future work.
@@ 3 CRATE-$\alpha$ Model
$\texttt{Prox}_{\eta,\lambda}[\bm{A};\bm{D},\bm{Z}]=\operatorname{ReLU}(\bm{A}-\eta\bm{D}^{\top}(\bm{D}\bm{A}-\bm{Z})-\eta\lambda\bm{1}).$ (9)
@@ 4 Experiments
Overall. The experimental section consists of three parts: (1) Scaling study: We thoroughly investigate the scaling behaviors of crate-$\alpha$ from Base to Large size and ultimately to Huge size. (2) Downstream applications: To further verify the broader benefits of scaling the crate-$\alpha$ model, we conduct additional experiments on real-world downstream tasks and present preliminary exploration results of crate-$\alpha$ on language tasks. (3) Interpretability: In addition to scalability, we study the interpretability of crate-$\alpha$ across different model sizes.
4.1 Dataset and Evaluation
@@
Scaling Study. For the transition from Base to Large size, we pre-train our model on ImageNet-21K and fine-tune it on ImageNet-1K via supervised learning. When scaling from Large to Huge, we utilize the DataComp1B [10] dataset within a vision-language pre-training paradigm, allowing us to study the effects of scaling the model to a massive size. For evaluation, we evaluate the zero-shot accuracy of these models on ImageNet-1K.
Downstream Applications. We include additional experimental results on four downstream datasets (CIFAR-10/100, Oxford Flowers, and Oxford-IIT Pets). We also examine the dense prediction capability of crate-$\alpha$ by training it on segmentation tasks using the ADE20K dataset [51]. For language tasks, we conduct new experiments with crate-$\alpha$ using autoregressive training on OpenWebText, following the setup in nanoGPT [16].
Interpretability. Following the evaluation setup of crate as outlined in [46], we apply MaskCut [43] to validate and evaluate the rich semantic information captured by our model in a zero-shot setting, including both qualitative and quantitative results.
4.2 Training and Fine-tuning Procedures
@@
Scaling Study. (1) Base to Large size: We initially pre-train the crate-$\alpha$ model on ImageNet-21K and subsequently fine-tune it on ImageNet-1K. During the pre-training phase, we set the learning rate to $8\times 10^{-4}$, weight decay to 0.1, and batch size to 4096. We apply data augmentation techniques such as Inception crop [35] resized to 224 and random horizontal flipping. In the fine-tuning phase, we adjust the base learning rate to $1.6\times 10^{-4}$, maintain weight decay at 0.1, and batch size at 4096. We apply label smoothing with a smoothing parameter of 0.1 and apply data augmentation methods including Inception crop, random horizontal flipping, and random augmentation with two transformations (magnitude of 9). For evaluation, we resize the smaller side of an image to 256 while maintaining the original aspect ratio and then crop the central portion to 224$\times$224. In both the pre-training and fine-tuning phases, we use the AdamW optimizer [24] and incorporate a warm-up strategy, characterized by a linear increase over 10 epochs. Both the pre-training and fine-tuning are conducted for a total of 91 epochs, utilizing a cosine decay schedule.
(2) Large to Huge size: In the pre-training stage, we utilize an image size of 84$\times$84, and the maximum token length is 32, with a total of 2.56 billion training samples. During the fine-tuning stage, we increase the image size to 224$\times$224 while maintaining the maximum token length at 32, with a 512 million training samples. Here, the key distinction between the pre-training stage and the fine-tuning stage is the image size. A smaller image size results in a faster training speed. In the configurations of crate-$\alpha$-CLIPA-B, crate-$\alpha$-CLIPA-L, and crate-$\alpha$-CLIPA-H, we use the crate-$\alpha$ model as the vision encoder, and utilize the same pre-trained huge transformer model from CLIPA [18] as the text encoder. For both the pre-training and fine-tuning stages, we freeze the text encoder and only train the vision encoder, i.e., the crate-$\alpha$ model. As we will show in the later results, this setup effectively demonstrates the scaling behaviors of crate-$\alpha$ models in the image domain. Detailed hyperparameter settings can be found in Appendix A.
Downstream Applications. On four downstream datasets, we follow the training setup from [46]. For the segmentation task, we compare the performance of CRATE and crate-$\alpha$ on the ADE20K dataset, mainly following the setup of [30] with minor modifications. Our batch size is set to 128, and the total number of training steps is 5000. For the language task, we conduct experiments with crate-$\alpha$ using autoregressive training on OpenWebText, following the setup in [16]. We compare crate-$\alpha$ models (57M and 120M) with CRATE and GPT-2, using results from CRATE reported in [45].
@@
Models (Base) IN-1K(%) Models (Large) IN-1K(%)
@@
4.3 Results and Analysis
Scaling the crate-$\alpha$ Model from Base to Large. As shown in Table 1, we compare crate-$\alpha$-B and crate-$\alpha$-L at patch sizes 32, 16, and 8. Firstly, we find our proposed crate-$\alpha$-L consistently achieves significant improvements across all patch sizes. Secondly, compared with the results of the vanilla crate (the first row of Table 1), increasing from crate-B to crate-L results in only a 0.5% improvement on ImageNet-1K. This indicates a case of diminishing returns. These findings compellingly highlight that the scalability of crate-$\alpha$ models significantly outperforms that of the vanilla crate. Meanwhile, the training loss in the pre-training stage is presented in Fig. 3; as the model capacity increases, the trend of the training loss improves predictably. This phenomenon is also described in [9].
Scaling the crate-$\alpha$ Model from Large to Huge. From the results shown in Fig. 4, we find that: (1) crate-$\alpha$-CLIPA-L/14 significantly outperforms crate-$\alpha$-CLIPA-B/16 by 11.3% and 9.0% in terms of ImageNet-1K zero-shot accuracy during the pre-training and fine-tuning stages, respectively. The substantial benefit suggests that the quality of learned representation may be constrained by the model size. Therefore, increasing the model size effectively leverages larger amounts of data. (2) When continuing to scale up model size, we also observe that crate-$\alpha$-CLIP-H/14 continues to benefit from larger training datasets, outperforming crate-$\alpha$-CLIP-L/14 by 3.1% and 2.5% in terms of ImageNet-1K zero-shot accuracy during the pre-training and fine-tuning stages, respectively. This demonstrates the strong scalability of the crate-$\alpha$ model. To explore the performance ceiling, we train a standard ViT-CLIPA-H/14 from scratch and observe improved performance.
Downstream Applications. On four downstream datasets, as shown in Table 2, we find that crate-$\alpha$ consistently outperforms CRATE, with both models pre-trained on IN21K, while crate-$\alpha$ demonstrates improved performance as model size increases. For the segmentation task, results in Table 3 show that crate-$\alpha$ consistently outperforms CRATE across all key metrics, with both models pre-trained on IN21K. These findings indicate significant performance gains in vision tasks beyond classification. For the language task, Table 4 shows that crate-$\alpha$ significantly improves over CRATE in language modeling. Due to limited time and resource constraints, we completed 80% of the total iterations for crate-$\alpha$-small and 55% for crate-$\alpha$-base, compared to the 600K total iterations used for CRATE. Nevertheless, crate-$\alpha$ still demonstrated notable improvements.
Interpretability. As shown in Fig. 5, we provide the segmentation visualization on COCO val2017 [20] for crate-$\alpha$, crate, and ViT, respectively. We find that our model preserves and even improves the (semantic) interpretability advantages of crate. Moreover, we summarize quantitative evaluation results on COCO val2017 in Table 6. Interestingly, when scaling up model size for crate-$\alpha$, the Large model improves over the Base model in terms of object detection and segmentation.
Table 2: The performance comparison between CRATE and CRATE-$\alpha$ across various datasets.
Dataset CRATE-B/32 CRATE-$\alpha$-B/32 CRATE-$\alpha$-L/32 CRATE-$\alpha$-B/16 CRATE-$\alpha$-L/14
CIFAR-10 97.22 98.17 98.68 98.67 99.10
CIFAR-100 85.27 89.40 91.16 90.58 92.57
Oxford Flowers-102 93.90 97.77 99.01 99.27 99.56
Oxford-IIIT-Pets 80.38 88.19 90.46 92.70 93.98
Table 3: Performance comparison of CRATE models with different configurations.
Model Scope mIoU mAcc aAcc
crate-$\alpha$-B/32 global 35.35 45.28 77.63
CRATE-B/32 global 30.28 39.29 75.21
Table 4: The comparison between CRATE and CRATE-$\alpha$ on the NLP task using the OpenWebText dataset.
GPT-2-base CRATE-base CRATE-$\alpha$-small CRATE-$\alpha$-base
Model size 124M 60M 57M 120M
CE val loss 2.85 3.37 3.28 3.14
@@ 4.4 Compute-efficient Scaling Strategy
We further explore methods to scale models efficiently in terms of computation. Table 1 demonstrates that the crate-$\alpha$ model scales effectively from the Base model to its larger variants. However, the pre-training computation for the top-performing model, crate-$\alpha$-L/8, is resource-intensive on ImageNet-21K. Inspired by CLIPA [18], we aim to reduce computational demands by using reduced image token sequence lengths, while maintaining the same training setup during the fine-tuning stage. The results are summarized in Table 5.
Table 5: Compute-efficient scaling strategy. To reduce the compute requirements of the pre-training stage, we use a model with a larger patch size. This results in a shorter token length for the same input size. The second and fourth columns indicate the compute requirements for the pre-training and fine-tuning stages, respectively, measured in TPU v3 core-hours. Details are provided in Section 4.4.
Pre-train Core-hours Fine-tune Core-hours Total core-hours IN-1K(%)
crate-$\alpha$-L/32 2,652 crate-$\alpha$-L/14 872 3,524 83.7
crate-$\alpha$-L/8 3,486 6,138 84.2
crate-$\alpha$-L/14 8,947 crate-$\alpha$-L/14 872 9,819 83.9
crate-$\alpha$-L/8 35,511 crate-$\alpha$-L/8 3,486 38,997 85.1
Results and analysis. (1) When fine-tuning with crate-$\alpha$-L/14 and using crate-$\alpha$-L/32 for pre-training on ImageNet-21K, this approach consumes about 35% of the TPU v3 core-hours required by crate-$\alpha$-L/14, yet achieves a promising 83.7% top-1 accuracy on ImageNet-1K, comparable to the 83.9% achieved by crate-$\alpha$-L/14; (2) When fine-tuning with crate-$\alpha$-L/8 and using crate-$\alpha$-L/32 for pre-training, this approach consumes just 15% of the training time required by crate-$\alpha$-L/8, yet it still achieves a promising 84.2% top-1 accuracy on ImageNet-1K, compared to 85.1% when using the crate-$\alpha$-L/8 model in the pre-training stage; (3) While the total computational cost of crate-$\alpha$-L/32 + crate-$\alpha$-L/8 is less than that of crate-$\alpha$-L/14 + crate-$\alpha$-L/14, the performance of the former is slightly better. In summary, we find that this strategy offers a valuable reference for efficiently scaling crate-$\alpha$ models in the future.
@@ 5 Discussion
Societal impact. A possible broader implication of this research is the energy consumption needed to conduct the experiments in our scaling study. However, there is growing interest in developing white-box transformers for better interpretability and transparency across a wide range of tasks and domains, including image segmentation [46], self-supervised masked autoencoders [26], and integrated sensing and communications [50], etc. Moreover, our results on the scalability of white-box transformers could also shed light on scaling up a broader class of white-box deep neural networks, such as white-box ISTA networks and their variants [11, 34, 3, 48, 17], designed via unrolled optimization. In summary, we believe that our findings and insights could be helpful for developing white-box transformers for a wide range of applications and tasks, benefiting a broad audience interested in building more interpretable and performant deep learning models and further amortizing the pre-training compute costs.
@@ Appendix A Additional Experiments and Details
A.1 Model configuration.
We provide details about crate-$\alpha$ model configurations in Table 7.
@@
Table 8: The comparison between CRATE-$\alpha$ and ViT. FLOPs and throughput are calculated based on an input size of 224x224 on an NVIDIA RTX A6000 graphics card.
Model FLOPs (G) #Params (M) Throughput Model FLOPs (G) #Params (M) Throughput
CRATE-$\alpha$-B/32 6.4 74.0 499 ViT-B/32 4.4 88.2 706
CRATE-$\alpha$-B/16 25.8 72.3 233 ViT-B/16 17.6 86.5 375
CRATE-$\alpha$-L/32 22.8 256.0 215 ViT-L/32 15.4 306.5 329
CRATE-$\alpha$-L/14 119.7 253.7 56 ViT-L/14 81.1 304.1 85
A.2 Comparison of model structure with ViT.
We also compare crate-$\alpha$ to ViT in terms of computational costs, number of parameters, and inference speed. These comparisons are summarized in Table 8, where crate-$\alpha$ matches ViT’s efficiency while achieving similar accuracy. With the same number of layers and embedding dimensions, crate-$\alpha$ has fewer parameters than ViT, and its FLOPs/Throughput is slightly higher.
To more accurately compare crate-$\alpha$ and ViT with larger model sizes, we conduct experiments on crate-$\alpha$-L/16 with an image resolution of 336, nearly matching the setup of ViT-L/16. Both models use a similar amount of FLOPs: 210G for crate-$\alpha$-L/16 compared to 191G for ViT-L/16. The throughput, or images processed per second, is also comparable at 35.53 for our model versus 35.56 for ViT-L/16. The accuracy of crate-$\alpha$-L/16 reach 84.6%, closely approaching ViT’s 85.2% under similar conditions. Meanwhile, combining the trend from Figure 1 (right) in the main paper, this narrowing performance gap from Base to Large model size suggests that crate-$\alpha$ can nearly matche ViT’s performance in large-scale settings. Besides, crate-$\alpha$ inherits the mathematical interpretability of the white-box models and can also achieve much better semantic interpretability evaluated by zero-shot segmentation.
A.3 Training details of crate-$\alpha$-CLIPA models.
When employing the crate-$\alpha$ architecture to replace the vision encoder in the CLIPA [18] framework, we essentially follow the original CLIPA training recipe. The setup for the pre-training stage is presented in Table 9. During the fine-tuning stage, we made some modifications: the input image size is set to $224\times 224$, the warmup steps are set to 800, and the base learning rate is set to 4e-7. When calculating the loss, we use the classification token from the vision encoder as the image feature and the last token from the text encoder as the text feature.
@@
Figure 6: One layer of the CRATE-$\alpha$ model architecture (with more details for the three modifications described in Section 3.
@@
Visualization of loss curves. We visualize the training loss curves of the four models, including crate and its three variants, in Fig. 8. We visualize the training loss curves of crate-$\alpha$-Base with different patch sizes in Fig. 9. In Fig. 10, we also visualize the training loss curves of models trained with efficient scaling strategy described in Section 4.4 in the main paper.
Figure 8: Training loss curves of different model architectures (mentioned in Fig. 1 in the main paper) on ImageNet-21K. The patch size is 32 for all four models shown in this figure. (+O: +overcomplete dictionary, +D: +decoupled dictionary, +R: +residual connection.)